OpenStudy (anonymous):

need help Match the probability with the outcome if two number cubes are rolled 100 times. an even number appeared on at least one cube 76 times

OpenStudy (anonymous):

anybody

OpenStudy (anonymous):

the probability an even number appears on "at least" one cube is \(\frac{3}{4}\) is that more or less clear?

OpenStudy (anonymous):

i need a percentage

OpenStudy (anonymous):

@satellite73 not clear

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

that is not what i am asking we can turn any number in to a percent, move the decimal place over two the question is, it is clear that on any one throw, the probability you get at least one even number is \(\frac{3}{4}\)

OpenStudy (anonymous):

so what are you saying

OpenStudy (anonymous):

@dlearner think of it like this: it is the same as flipping a coin, half even, half odd (e, e, ) (e, o), (o, e) , (o,o) three out of four times you get at least one even

OpenStudy (anonymous):

what would it be in percentage then ????

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

you are not close to done now you need to use the binomial with \(n=100, k=76, p=\frac{3}{4},1-p=\frac{1}{4}\)

OpenStudy (anonymous):

then figure out the percentage by moving the decimal that is not really much of the problem